In this assignment you will be working with bonds. You will compute bond issue price, then make an amortization table. You will then do a few select journal entries using your amortization table.

Here is the data:

  • Bond issue principal amount: 800,000
  • Semi-annual interest payments at the end of every six months
  • Term = 6 years
  • Stated rate = 6%
  • Interest payment = ?

Scenarios:

  1. Market rate = 4%
  2. Market rate = 6%
  3. Market rate = 8%

Requirements:

For each of the three scenarios, complete a separate tab in Excel. Include the data for your problem in each sheet. Then:

  1. Compute the interest payment. It is the same amount in each scenario.
  2. Compute bond issue price
  3. Create an amortization table using the effective interest method for all 12 payment periods.
  4. Below your Effective Interest amortization table show journal entries for:
    1. Bond Issue
    2. First payment (payment at end of period 1)
    3. Payment at the end of the 6th period.
    4. Settlement of the bond after last payment. Note: do not include the final payment journal entry; only include the entry for the settlement of bon
  5. For the instances that have a discount or a premium (scenarios 1 and 3) do the following:
    1. Below your journal entries, create an amortization table using the same bond issue price as you did in step 2 above. Use the straight-line method of amortization for this second table (see chapter 14 appendix).
  6. Below your straight-line amortization table show journal entries for:
    1. First payment (payment at end of period 1)
    2. Payment at the end of the 6th period.
Answer #1
Interest payment=Principal amount*Interest rate for 6 months=800000*6%*(6/12)=$ 24000
Market rate = 4%:
Issue price of the bond=Present value of interest payments+Present value of maturity value at the end
Number of semi-annual periods=6*2=12
Discount rate=Bonds yield for 6 months=4%*(6/12)=2%
Present value of interest payments=Interest payment*PV factor at 2% for 12 years=24000*10.57534=$ 253808
Present value of maturity value at the end=Maturity value*PV factor at 2% for 12th year=800000*0.78849=$ 630792
Issue price of the bond=253808+630792=$ 884600
Premium on issue of bonds=Issue price-Maturity value=884600-800000=$ 84600
Bond amortization table
Date Cash paid Interest expense Premium amortized Carrying value of the bond
a b a-b
Beg. 884600
End of 1st period 24000 17692 6308 878292
End of 2nd period 24000 17566 6434 871858
End of 3rd period 24000 17437 6563 865295
End of 4th period 24000 17306 6694 858601
End of 5th period 24000 17172 6828 851773
End of 6th period 24000 17035 6965 844808
End of 7th period 24000 16896 7104 837705
End of 8th period 24000 16754 7246 830459
End of 9th period 24000 16609 7391 823068
End of 10th period 24000 16461 7539 815529
End of 11th period 24000 16311 7689 807840
End of 12th period 24000 16160 7840 800000
Cash paid=Interest payment
Interest expense=Beginning carrying value of the bond*Bonds yield for 6 months (2%)
Carrying value of the bond=Beginning carrying value of the bond-Premium amortized
Journal entry:
Ref. Account titles Debit Credit
a Cash 884600
Bonds payable 800000
Premium on bonds payable 84600
(Bond Issue)
b. Interest expense 17692
Premium on bonds payable 6308
Cash 24000
(payment at end of period 1)
c. Interest expense 17035
Premium on bonds payable 6965
Cash 24000
(Payment at the end of the 6th period)
d. Interest expense 16160
Premium on bonds payable 7840
Cash 24000
(Payment at the end of the life of bond)
Bonds payable 800000
Cash 800000
(Settlement of the bond)

Market rate = 6% Issue price of the bond+Present value of interest payments+Present value of maturity value at the end Number24000 24000 0 800000 End of 10th period End of 11th period End of 12th period 0 24000 24000 24000 24000 800000 800000 0 Cash800000 Cash (Settlement of the bond) Market rate = 8%: Issue price of the bond=Present value of interest payments+Present val29846 5846 751987 6079 758066 24000 24000 24000 24000 6323 End of 5th period End of 6th period End of 7th period End of 8th pc. 30079 6079 24000 d. 31688 (payment at end of period 1) Interest expense Discount on bonds payable Cash (Payment at the end